Solve the system of equations.
{-x + y + 2z = 0
{ x + 2y + x = 6
{-2x - y + z = -6
Solve for x, y and z.

Answer by TimothyLamb(4379)

---
-x + y + 2z = 0
x + 2y + x = 6
-2x - y + z = -6
---
put the system of linear equations into standard form
---
-x + y + 2z = 0
2x + 2y = 6
-2x - y + z = -6
---
copy and paste the above standard form linear equations in to this solver:
https://sooeet.com/math/system-of-linear-equations-solver.php
---
answer:
this linear system is singular, it has no solution
---
Solve and graph linear equations:
https://sooeet.com/math/linear-equation-solver.php
---
Solve quadratic equations, quadratic formula:
https://sooeet.com/math/quadratic-formula-solver.php
---
Solve systems of linear equations up to 6-equations 6-variables:
https://sooeet.com/math/system-of-linear-equations-solver.php